HUBBLE SPACE TELESCOPE DAILY REPORT #823 PERIOD COVERED: 0800 EST Daily Status Report as of 046/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: read 13 1.1 Completed FOS/BL 3791 (Quasar Absorption Line Survey) The Faint Object Spectrograph (Red Side) attempted to make observations of quasar PKS1302-10. The baseline guide star acquisition failed. All seven observations were reported as having failed. (See Operations Note 2.4) 1.2 Completed Two Sets of FOC/96 4107 (Search QSOs Suitable for Subsequent Observations) The Faint Object Camera attempted to make two separate observations of quasar 1320-106. Both baseline guide star acquisitions failed after FGS #2 electronics had been turned off. The shutter never opened for either observation. (See Operations Note 2.4) 1.3 Completed FOC/96 3504 (High Spatial Resolution Imaging and Spatial of Active Galactic Nuclei) The Faint Object Camera was used to attempt to make observations of active galactic nuclei 3C277.3 . The base line guide star acquisition failed because the FGS #2 electronics had been turned off. The shutter never opened. 1.4 Completed WFC 1297 (GTO High Latitude Parallel Program) The Wide Field Camera attempted to make two images which were to have been taken in parallel with the prime science observations. However, the baseline guide star acquisition on the prime science observation failed, due to the FGS #2 electronic having been turned off. The shutter never opened. (See Operations Note 2.4) 1.5 Completed Two Sets of FGS 3688 (Diagnostic of Early Stellar Evolution) Two sets of observations were scheduled using the FGS to make astrometric observations of pre-main sequence star 100628. These astrometric observations did not occur, because the guide star acquisition failed due to the ongoing problems with FGS #2. (See Operations Note 2.4 ) 1.6 Uplinked Updated Ephemeris Tables, Per MOSES 2871 1.7 Completed WFC-PC 4109 (Rapid Internal Monitor) Both the Wide Field Camera and the Planetary Camera were used to make two internal flat observations using filter F555W to check for new forms of contamination (i.e., Daisies and measles). Both observations were received and displayed. The two internal flats were reported to look normal. 1.8 Completed GHRS/2 3755 (Absorption Cross Sections of Nearby Galaxies) The High Resolution Spectrograph was used to attempt to make observations of quasar 1543+489 to study the absorption lines produced by nearby galaxies. The science observations were not executed due to the problems with FGS #2. The wavelenght calibration was also not taken. 1.9 Completed FOS/RD 4059 (Absolute Photometric Monitor) The Faint Object Spectrograph was used to make eight observations of standard star HZ-44 to help determine the absolute sensitivity of the FOS detectors. These observations will be used for pipeline data calibrations. All eight observations were received and displayed. The guide star acquisitions did no occur. The spectral observations contained some light, the target was reported to not be properly centered. 1.10 Completed Three Sets of PC 3603 (Snapshot of Protoplanetary Nebulae) The Planetary Camera was used to make three sets of observations of the protoplanetary nebula CPD-64D293. The first observations did occur in spite of the fact that the guide star acquisition did not occur. (See operations Note 2.4) Its image was smeared. Of the remaining subsequent images; one showed a faint stellar object, one looked normal, and three were reported to be blank. 1.11 Completed GHRS/2 2537 (Physical Conditions of Interstellar Gas) The High Resolution Spectrograph was used to attempt to obtain high signal to noise spectra of star HD80007. The deflection calibration observation occurred. The remaining four science observations did not occur because of problems using FGS #2. (See Operations Note 2.4) 1.12 Completed PC 4249 (Black Holes Stellar Dynamics and Populations) The Planetary Camera was used to take two observations of the inner core region of the elliptical galaxy NGC-4623 which is located in the Virgo Cluster. Both observations failed because the FGS #2 was turned off. The shutter toggled open for 0.11 seconds, but nothing was noticed in the field of view. 1.13 Completed GHRS/2 4065 (Dark Noise Monitoring) The High Resolution Spectrograph was used to make one internal dark calibration observation. The observations was received and displayed in the OSS. It was reported to look normal. 1.14 Completed FOS/RD 4698 (FOS Dark Monitoring) The Faint Object Spectrograph was used on both the red and blue sides to obtain ten dark exposures to determine the instruments background calibration. All ten observations were received and displayed. All ten were reported to look normal. 1.16 Initiate GHRS/2 4766 (Boron in Population II Dwarfs, Part 2) The High resolution Spectrograph was used to initiate spectral observations of the Boron I lines in the metal poor unevolved star HD140283. The OSS monitored the spiral search portion of the onboard target acquisition. The target was found in the first map. The observations are currently ongoing, and will be analyzed later. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 18 Successful : 2 (16 failures due to FGS #2, taken off line at 043/1829Z Scheduled Reacquisitions : 26 Successful : 8 (18 failures due to FGS #2 problems) Losses of Lock : 0 2.2 FHST Updates: Scheduled : 35 Successful : 35 2.3 Significant HSTARs: None 2.4 Operations Note: FGS #2 has shown an anomalous behavior where it intermittently fails to move to certain positions. The problem is yet undetermined, and data continues to be analyzed at this time. A special operations request executed to shut down FGS #2 at 043/1435Z. This will make the unit thermally stable, but will not allow it to accept normal operational commanding. The STScI is regenerating SMSs to restrict FGS #2 usage. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete GHRS/2 4766 (Boron in Population Two Dwarf Stars, Part 2) 3.2 Complete FOC/96 3325 (Structure of Globular Clusters) 3.3 Complete FOC/96 4353 (Far UV Observations of the Giant Planets) 3.4 Complete FOS/BL 4698 (FOS Dark Monitoring) 3.5 Initiate FOC/(^ 4071 (UV Studies of Galaxies) 3.6 Initiate WFC 1297 (GTO High Latitude Parallel Program) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. -END- ----- End Included Message -----